# GTPO and GRPO-S: Token and Sequence-Level Reward Shaping with Policy Entropy

### v5

@@ 2.3 A Sequence-Level Variant of GTPO (GRPO-S)
Sequence-Level Reward Shaping. For any sequence $o_{k}\in O$, rewards are shaped based on a sequence’s average token entropy, $\hat{H}_{k}=\frac{1}{|o_{k}|}\sum_{t=1}^{|o_{k}|}H_{k,t}$. For successful sequences ($o_{i}\in O^{+}$), the reward is augmented with an entropy-based bonus to reinforce valuable exploration. Conversely, for unsuccessful sequences ($o_{j}\in O^{-}$), an additional penalty proportional is applied to their average inverse entropy, thus penalizing high-confidence mistakes more severely. Formally:
$\hat{r}_{i}^{+}=\beta_{1}r_{i}+\beta_{2}\frac{\hat{H}_{i}}{\sum_{k=1}^{n}\hat{H}_{k}}\cdot n\qquad\text{and}\qquad\hat{r}_{j}^{-}=\beta_{1}\cdot(-1)+\beta_{2}\frac{1/\hat{H}_{j}}{\sum_{k=1}^{m}(1/\hat{H}_{k})}\cdot m\cdot(-1),$ (6)
where $\beta_{1},\beta_{2}>0$ are hyperparameters. This formulation rewards successful sequences that are, on average, more exploratory, while penalizing confidently incorrect sequences.
The GRPO-S Objective Function. The advantage functions $\hat{A}_{i}^{+}$ and $\hat{A}_{j}^{-}$ are computed analogously to Equation 1, but using the sequence-level shaped rewards and normalizing over the $G$ sequences in the group. The final objective function for GRPO-S mirrors the structure of the original GRPO loss, but with our shaped advantages:
$\begin{split}\mathcal{J}_{\text{GRPO-S}}(\theta)=\mathbb{E}\bigg[\frac{1}{G}\bigg(&\sum_{i=1}^{n}\min\big(\hat{w}_{i}(\theta)\hat{A}_{i}^{+},\text{clip}(\hat{w}_{i}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i}^{+}\big)\\ +&\sum_{j=n+1}^{G}\min\big(\hat{w}_{j}(\theta)\hat{A}_{j}^{-},\text{clip}(\hat{w}_{j}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{j}^{-}\big)\bigg)\bigg],\end{split}$ (7)
where the sequence-level importance weight $\hat{w}_{i}(\theta)$ averages the token-level weights:
@@ 2.4.1 Implementation Details
Robust Concept Definitions via Geometric Mean. The arithmetic mean for aggregating sequence-level importance sampling weights can cause training instability. The underlying ratio-based weights have a skewed distribution, making the arithmetic mean sensitive to outliers. To mitigate this, the robust geometric mean is employed. By averaging in logarithmic space, the geometric mean is better suited for ratios and dampens the influence of extreme values, yielding a stable aggregation. The sequence-level importance weight is thus redefined as follows:
$\tilde{r}_{i,t}$ $=\alpha_{1}r_{i}+\alpha_{2}\frac{H_{i,t}}{(\prod_{k=1}^{|o_{i}|}H_{k,t})^{1/|o_{i}|}},$ $\tilde{r}_{j,t}^{-}=\alpha_{1}(-1)+\alpha_{2}\frac{1/H_{j,t}}{(\prod_{k=1}^{|o_{j}|}1/H_{k,t})^{1/|o_{j}|}}(-1),$
$\hat{r}_{i}^{+}$ $=\beta_{1}r_{i}+\beta_{2}\frac{\hat{H}_{i}}{(\prod_{k=1}^{n}\hat{H}_{k})^{1/n}},$ $\hat{r}_{j}^{-}=\beta_{1}(-1)+\beta_{2}\frac{1/\hat{H}_{j}}{(\prod_{k=1}^{|o_{j}|}1/\hat{H}_{k})^{1/|o_{j}|}}(-1),$
@@
Initialize: Policy parameters $\theta$, reference policy $\pi_{\theta_{\text{ref}}}$, hyperparameters $\alpha,\beta,G,\epsilon$.
for each training iteration $k=1,2,\dots$ do
  Sample a batch of prompts $\{q\}$.
  $\pi_{\theta_{old}}\leftarrow\pi_{\theta}$
  For each prompt $q$, sample a group of $G$ sequences $\{o_{1},\dots,o_{G}\}$ using $\pi_{\theta_{old}}$.
  For each sequence $o_{i}$, compute its terminal reward $r_{i}\in\{0,1\}$.
  Partition the group into successful sequences $O^{+}$ and unsuccessful $O^{-}$.
  Compute policy entropy $H_{i,t}$ for each token $o_{i,t}$ using $\pi_{\theta_{old}}$.
  — GTPO Branch (Token-Level) —
  For each token $o_{i,t}\in O^{+}$ and $o_{j,t}\in O^{-}$, compute $\tilde{r}_{i,t}^{+}$ and $\tilde{r}_{j,t}^{-}$ using Eq. (2) and (3).
  Compute token-level advantages $\tilde{A}_{i,t}^{+}$ and $\tilde{A}_{j,t}^{-}$ using Eq. (4).
  Compute the GTPO loss $\mathcal{J}_{\text{GTPO}}(\theta)$ using Eq. (5).
  Update $\theta$ using a gradient step on $\mathcal{J}_{\text{GTPO}}(\theta)$.
  — GRPO-S Branch (Sequence-Level) —
  For each sequence $o_{i}\in O^{+}$ and $o_{j}\in O^{-}$, compute average entropy $\hat{H}_{i}$ and $\hat{H}_{j}$.
  Compute shaped sequence rewards $\hat{r}_{i}^{+}$ and $\hat{r}_{j}^{-}$ using Eq. (6).
  Compute sequence-level advantages $\hat{A}_{i}^{+}$ and $\hat{A}_{j}^{-}$.
  For each sequence, compute the geometric importance weight $\hat{w}_{i}(\theta)$ using Eq. (8).
  Compute the GRPO-S loss $\mathcal{J}_{\text{GRPO-S}}(\theta)$ using Eq. (7).
  Update $\theta$ using a gradient step on $\mathcal{J}_{\text{GRPO-S}}(\theta)$.
end for
@@ 2.4.2 Theoretical Guarantees
This section provides a theoretical analysis establishing that our proposed reward shaping mechanisms preserve the expected policy gradient direction of the baseline algorithms, a key condition for ensuring convergence. Our approach redistributes rewards to create a more granular, token-aware learning signal. As the entropy term is detached from the gradient computation, our analysis demonstrates that while these modifications alter training dynamics, they maintain the expected gradient direction, thus guiding optimization towards a valid policy optimum.
Analysis of the Token-Level Objective (GTPO). Our token-level objective modifies the reward structure of the GRPO baseline. As our reward shaping applies exclusively to successful sequences (where the original reward $r_{i}=1$), the analysis centers on this redistribution. By construction, the shaping is conservative; the weighted redistribution of token-level rewards $\tilde{r}_{i,t}^{+}$ for these positive sequences preserves the total reward:
$\sum_{i,t}\tilde{r}_{i,t}^{+}=(\alpha_{1}+\alpha_{2})\sum_{i=1}^{G}|o_{i}|r_{i}\approx\sum_{i=1}^{G}|o_{i}|r_{i}.$
The equations hold because $\alpha_{1}+\alpha_{2}\approx 1$ has been set. The conservation of total reward leads directly to the conclusion that the expected mean reward remains unchanged:
$\mathbb{E}\big[\text{mean}\big(\{\hat{r}_{k,t}^{+}\}_{1\leq k\leq G,\ 1\leq t\leq|o_{k}|}\big)\big]\approx\mathbb{E}\big[\text{mean}\big(\{r_{k,t}\}_{1\leq k\leq G,\ 1\leq t\leq|o_{k}|}\big)\big],$
where $\text{mean}({R})=\frac{\sum_{i}|o_{i}|r_{i}}{\sum_{i}|o_{i}|}$. This equivalence, in turn, implies that the expectation of the advantage function under our proposed reward shaping, $\mathbb{E}[\tilde{A}_{i,t}^{+}]$, is approximately equal to the expectation of the baseline advantage function, $\mathbb{E}[\hat{A}_{i,t}]$ ( $\mathbb{E}[\tilde{A}_{i,t}^{+}]\approx\mathbb{E}[\hat{A}_{i,t}]$), where $\hat{A}_{i,t}=\frac{r_{i}-\text{mean}({R})}{\text{std}({R})}$. The source of this approximation and a detailed description of $\text{mean}({R})$ are provided in Appendix B.4.
Since the entropy term is detached, the expected policy gradient $\mathbb{E}$ approximates that of the GRPO baseline ($\mathbb{E}[\nabla_{\theta}\mathcal{J}_{\text{GTPO}}(\theta)]\approx\mathbb{E}[\nabla_{\theta}\mathcal{J}_{\text{DAPO}}(\theta)]$). This preserves the expected gradient direction, ensuring optimization towards the same local optimum. The modification primarily impacts training dynamics by altering the gradient estimator’s variance. This redistribution provides a fine-grained, token-level signal that can yield a lower-variance estimator, potentially stabilizing and accelerating convergence, as detailed in Appendix B.2. A parallel analysis for our sequence-level objective (GRPO-S), which similarly preserves the expected policy gradient while aiming to reduce estimator variance, is provided in Appendix B.3.
3 Experiments
3.1 Experimental Setup
Tasks and Datasets. Methods are evaluated on the AIME 2024 and AIME 2025 benchmarks. These challenging mathematical datasets require long-horizon, chain-of-thought reasoning, making them a rigorous testbed for assessing advanced alignment techniques.
Evaluation Metrics. Our primary metric is Pass@k ($k\in\{2,4,8,16,32\}$), which encourages solution diversity over the more conservative Pass@1. mean32 is also reported. Improvements are quantified by Absolute (APG) and Relative (RPG) Performance Gains over the baseline.
Models and Baselines. Experiments are conducted on Qwen2.5-7B and Qwen2.5-32B models. Comparisons are made against faithful implementations of GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025), both state-of-the-art alignment baselines, to ensure a strong comparison.
Implementation Details. Experiments were run on 64 GPUs with a global batch size of 128, a group size of 16, and a learning rate of $1\times 10^{-6}$. For generation, a temperature 1.0 and top-p 1.0 are used, with max lengths of 2048 (prompt) and 4096 (response). Key reward shaping hyperparameters are $\alpha_{1}=\beta_{1}=1$, $\alpha_{2}=\beta_{2}=0.1$, and entropy is clipped at $\epsilon_{low}=0.2,\epsilon_{high}=0.28$.
@@ 3.1 Experimental Setup
[…]
Table 1: Performance of GTPO and GRPO-S against the DAPO and GRPO baselines on AIME 2024 and 2025 benchmarks, reporting maximum Pass@k and mean32 scores. APG/RPG denote Absolute/Relative Performance Gain over DAPO. Our methods show significant, consistent improvements. *Pass@k on AIME 2025 plateaus for k ¿ 4 due to the limited test set size.
3.2 Comparative Performance Analysis
As presented in Table 1, our methods, GTPO and GRPO-S, establish a new state-of-the-art by consistently and substantially outperforming both the GRPO and DAPO baselines across all configurations. Our approach critically resolves a trade-off observed in prior methods: while a strong baseline like DAPO demonstrates competitive performance, its stability-focused loss can limit exploration, causing it to fall short on Pass@k metrics in certain settings (e.g., AIME 2025). In stark contrast, our entropy-based reward shaping successfully elevates both the overall performance metric (mean32) and these exploration-sensitive scores. The efficacy of this mechanism is particularly pronounced on smaller models; for instance, GTPO’s relative performance gain (RPG) over DAPO on the AIME 2025 benchmark reaches a remarkable +60.0% on the 7B model, compared to +37.5% on the 32B model. This suggests our entropy-weighting provides a crucial learning signal for smaller models, which are more susceptible to premature convergence, by encouraging exploration and helping them navigate complex reasoning spaces more effectively.
3.3 Reward Trajectories and Sample Efficiency
The mean reward trajectories on the test set (Fig. 4) illuminate the benefits of our approach, demonstrating that our methods not only achieve a significantly higher final reward ceiling but do so with strong sample efficiency. The models largely converge within 210 training steps, a finding substantiated by the training set reward curves, which are deferred to Appendix D.4 for brevity. This rapid convergence indicates that the substantial benefits of enhanced exploration do not come at the cost of slower learning. Across all datasets and model sizes, GTPO consistently reaches the highest reward plateau, followed closely by GRPO-S, with both substantially outperforming DAPO.
Our methods foster sustained exploration and prevent policy collapse, establishing a clear causal chain from our entropy-weighted reward to superior performance. A detailed analysis and the empirical evidence for this mechanism are presented in Appendix D.1.
3.4 Hyperparameter Sensitivity Analysis
Figure 3: Hyperparameter comparison.
We conducted a sensitivity analysis on the key reward shaping hyperparameters for GTPO and GRPO-S, with results presented in Fig. 3. Across all tested configurations, both of our methods demonstrate robust and significant performance gains over the DAPO baseline on mean32 and Pass@32 metrics. Among them, GRPO-S exhibits higher stability across settings, providing a clear view of the performance dynamics. As the weight of the entropy bonus ($\beta_{2}$) increases from 0.1 to 0.2, we observe a clear decline in overall performance for GRPO-S. This finding confirms that while exploration is beneficial, an excessive entropy bonus can detract from optimizing the primary task objective.
@@ 3.4 Hyperparameter Sensitivity Analysis
Figure 4: Mean reward trajectories on the test sets. All curves are smoothed for visual clarity. Each row corresponds to a different experimental setting: (Top) AIME 2024 with Qwen2.5-32B, (Middle) AIME 2025 with Qwen2.5-32B, and (Bottom) AIME 2025 with Qwen2.5-7B. Columns show different metrics from left to right: Mean Reward, Mean Reward of Pass@2, Pass@8, and Pass@32. For brevity and to maintain visual clarity, the corresponding results for Pass@4 and Pass@16, which exhibit similar trends, are deferred to Appendix D.3, see Fig. 7 and Fig. 8.
4 Discussion
Future Work. The concept of relative entropy itself warrants deeper exploration. While this work highlights its importance, the optimal method for comparison remains an open question. Three distinct approaches for measuring this relativity are identified. Visualizing the token entropies as a matrix $H$, where $H_{i,t}$ is the entropy of the $t$-th token in the $i$-th response, these methods can be understood as: a column-wise comparison, which compares the entropy of tokens at the same position across different responses (i.e., within a column of $H$); a row-wise comparison, which compares tokens at different positions within the same response (i.e., within a row of $H$); and a matrix-wise comparison, which compares all tokens from all responses collectively (i.e., across the entire matrix $H$). Further experimentation is needed to verify which of these approaches is superior. The conceptual difference is illustrated below:

### v6

@@ 2.3 A Sequence-Level Variant of GTPO (GRPO-S)
Sequence-Level Reward Shaping. For any sequence $o_{k}\in O$, rewards are shaped based on a sequence’s average token entropy,
$\hat{H}_{k}=\frac{1}{|o_{k}|}\sum_{t=1}^{|o_{k}|}H_{k,t}.$ (8)
For any successful sequence $o_{i}\in O^{+}$, the reward is augmented with an entropy-based bonus to reinforce valuable exploration. Conversely, for any unsuccessful sequence $o_{j}\in O^{-}$, an additional penalty proportional is applied to its average inverse entropy, thus penalizing high-confidence mistakes more severely. Formally:
$\begin{split}\hat{r}_{i}^{+}&=\beta_{1}r_{i}+\beta_{2}\frac{\hat{H}_{i}}{\sum_{k=1}^{n}\hat{H}_{k}}\cdot n,\\ \hat{r}_{j}^{-}&=\beta_{1}\cdot(-1)+\beta_{2}\frac{1/\hat{H}_{j}}{\sum_{k=1}^{m}(1/\hat{H}_{n+k})}\cdot m\cdot(-1).\end{split}$ (9)
where $\beta_{1},\beta_{2}>0$ are hyperparameters and $\hat{r}_{i}^{-}=\hat{r}_{j}^{+}=0$. This formulation rewards successful sequences that are, on average, more exploratory, while penalizing confidently incorrect sequences.
The GRPO-S Objective Function. The advantage functions $\hat{A}_{i}^{+}$ and $\hat{A}_{j}^{-}$ are computed analogously to Eq. (1), but using the sequence-level shaped rewards and normalizing over the $G$ sequences in the group. The final objective function for GRPO-S mirrors the structure of the original GRPO loss, but with our shaped advantages:
$\begin{aligned} \mathcal{J}_{\text{GRPO-S}}(\theta)=\mathbb{E}\bigg[\frac{1}{G}\bigg(&\sum_{i=1}^{n}\min\big(\hat{w}_{i}(\theta)\hat{A}_{i}^{+},\text{clip}(\hat{w}_{i}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i}^{+}\big)\\ +&\sum_{j=n+1}^{G}\min\big(\hat{w}_{j}(\theta)\hat{A}_{j}^{-},\text{clip}(\hat{w}_{j}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{j}^{-}\big)\bigg)\bigg],\end{aligned}$
where $\hat{w}_{i}(\theta)$ is the sequence-level importance weight that averages the token-level weights:
@@
Remark 2.1.
Note that the definitions of our new algorithms involve multiple fractional terms. In practical implementation, a small positive constant $\epsilon$ needs to be added where necessary to ensure numerical stability. For the sake of brevity and to avoid cluttering the definitions, we omit explicit notation of this constant in the text.
Table 1: Comparison of different methods on various benchmarks.
AIME 2024 AIME 2025 MATH 500
Method Mean@32 Pass@4 Pass@32 Mean@32 Pass@4 Pass@32 Mean@32
Qwen2.5-7B
GRPO 28.36 36.15 46.01 15.00 19.03 20.00 78.62
DAPO 21.2 30.04 35.51 13.33 16.28 16.67 78.85
DAPO w/ Forking Tokens 30.77 41.34 54.16 15.21 19.82 21.67 79.25
GTPO 34.23 48.03 64.89 16.67 25.51 26.67 80.16
GRPO-S 34.67 47.18 64.76 18.33 22.76 23.33 80.07
Qwen2.5-32B
GRPO 29.17 37.18 46.30 17.71 25.95 28.85 84.16
DAPO 34.06 46.64 59.02 21.67 25.76 26.67 84.45
DAPO w/ Forking Tokens 34.68 47.81 64.12 23.19 27.99 32.02 85.21
GTPO 35.21 49.41 68.91 26.89 33.49 36.67 85.72
GRPO-S 35.52 51.12 67.19 25.11 29.75 34.33 85.79
@@ 2.4.1 Implementation Details
0: Initial policy model $\pi_{\theta}$; Task prompts $\mathcal{D}$; Hyperparameters $\alpha,\beta,G,\epsilon$; Learning rate $\eta$
1:Initialize: $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
2:for iteration $k=1,2,\dots$ do
2:   Step 1: Data Collection & Entropy Calculation
3:  $\{o_{1},\dots,o_{G}\}\sim\pi_{\theta_{\text{old}}}(\cdot|q)$ # Sample $G$ outputs
4:  $r_{i}\leftarrow R(o_{i})\in\{0,1\},\forall i$ # Get terminal rewards
5:  $O^{+},O^{-}\leftarrow\text{Partition}(O)$ # Split based on outcome
6:  $H_{i,t}\leftarrow-\sum\pi_{\theta_{\text{old}}}\log\pi_{\theta_{\text{old}}}$ # Token-level entropy
6:   Step 2: Reward Shaping & Optimization
7:  if Method is GTPO (Token-Level) then
[…]
9:   $\tilde{A}_{*,t}^{+,-}\leftarrow\tilde{r}_{*,t}^{+,-}$ # Token Advantage (Eq. 6)
10:   $\mathcal{L}\leftarrow\mathcal{J}_{\text{GTPO}}(\theta;\tilde{A}_{*,t}^{+,-})$ # Loss via Eq. (7)
11:  else if Method is GRPO-S (Sequence-Level) then
12:   $\hat{H}_{i}\leftarrow\text{Avg}(H_{i,t})$ # Sequence entropy
[…]
14:   $\hat{A}_{*}^{+,-}\leftarrow\hat{r}_{*}^{+,-}$ # Sequence Advantage
[…]
16:   $\mathcal{L}\leftarrow\mathcal{J}_{\text{GRPO-S}}(\theta;\hat{A}_{i},\hat{w}_{i})$ # Loss via Eq. (10)
17:  end if
18:  $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{L}$ # Gradient Update
19:  $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$ # Sync policy
20:end for
[…]
@@ 2.4.2 Theoretical Guarantees
We provide a theoretical analysis establishing the unbiasedness and convergence of our framework, using GTPO as a representative case. Detailed proofs and a parallel analysis for GRPO-S are deferred to Appendices B.3 and B.4.
Proposition 2.2 (Conservation of Positive Reward Mass - GTPO).
@@
$\sum_{i\in O^{+}_{t}}\tilde{r}^{+}_{i,t}=\sum_{i\in O^{+}_{t}}r_{i}.$ (12)
Proposition 2.3 (Asymptotic Consistency - GTPO).
Let $\Delta\mathcal{J}(\theta)=\nabla\mathcal{J}_{\text{GTPO}}^{+}(\theta)-\nabla\mathcal{J}_{\text{GRPO}}(\theta)$ be the gradient bias introduced by the entropy-weighted reward shaping, where $\mathcal{J}_{\text{GTPO}}^{+}(\theta)$ is the positive reward component of GTPO. Assume the policy $\pi_{\theta}$ satisfies the Entropy Consolidation Condition: as training iteration $k\to\infty$, the variation in token entropy among successful sequences diminishes, i.e., for any two successful sequences $o_{i_{1}},o_{i_{2}}\in O^{+}$, $\lim_{k\to\infty}\frac{H_{{i_{1}},t}}{H_{{i_{2}},t}}=1$ (almost surely, assuming regularization $\epsilon>0$ prevents singularity). Then, the gradient bias vanishes asymptotically:
$\lim_{k\to\infty}\|\Delta\mathcal{J}(\theta_{k})\|=0.$ (13)
Theorem 2.4 (The Same Global Optimum - GTPO).
GTPO shares the same global optimum as DAPO.
Proof.
Proposition 12 establishes that the positive reward mass of GTPO is identical to that of GRPO. It is worth noting that, from a mathematical perspective, the design of DAPO is nearly identical to GRPO, with the primary distinction being the assignment of a reward of -1 to incorrect answers rather than the 0 reward used in GRPO. Consequently, following a derivation analogous to Proposition 2.1, we can conclude that the total reward mass of GTPO is equivalent to that of DAPO. The conservation of reward mass implies that the GTPO objective is an unbiased augmentation of the DAPO objective in terms of the total learning signal.
Note that the gradient estimator for the positive reward component of GTPO can be conceptually decomposed into:
$\mathbb{E}[\nabla\mathcal{J}_{\text{GTPO}}^{+}]=\mathbb{E}[\nabla\mathcal{J}_{\text{GRPO}}]+\underbrace{\mathbb{E}\left[\text{Cov}\left(\Delta_{\text{entropy}},\nabla\log\pi_{\theta}\right)\right]}_{\text{Exploration Term}},$ (14)
where $\Delta_{\text{entropy}}$ represents the zero-sum redistribution of rewards based on uncertainty. Since the total positive reward remains tied to the ground truth $r_{i}$, the model cannot maximize the objective solely by manipulating entropy; it must eventually increase the likelihood of correct answers to gain more total reward mass. Furthermore, as $k\to\infty$}, the entropy term $H_{i_{1},t},H_{i_{2},t}\to 0$ and become uniform across successful paths, causing $\frac{H_{{i_{1}},t}}{H_{{i_{2}},t}}\to 1$ and $\tilde{r}_{i,t}^{+}\to r_{i}$. Based on Proposition 13, we know that the positive reward component of GTPO shares the same global optimum as GRPO. By extension, given that DAPO corresponds to GRPO subject to a simple reward shift for incorrect answers that preserves the optimal policy, which is a modification consistent with our own reward design for incorrect answers, it follows that GTPO shares the same global optimum as DAPO. ∎
Remark 2.5.
Our analysis focuses on the reward redistribution mechanism, which generates a granular, token-aware learning signal. This shaping is sign-consistent with the original objective, amplifying positive outcomes while penalizing negative ones. Furthermore, as the entropy term is detached from the gradient computation, it reshapes the reward landscape to favor exploration. Consequently, our analysis confirms that while these modifications influence training dynamics, they ensure optimization towards a valid policy optimum.
Remark 2.6.
Note that while Proposition 12 assumes $\alpha_{1}+\alpha_{2}=1$, this condition is not strictly necessary. The proofs remain valid even if $\alpha_{1}+\alpha_{2}\neq 1$, as the term $\alpha_{1}+\alpha_{2}$ merely acts as a constant scaling factor that does not affect the validity of the derivation. In our experimental setup, the sum $\alpha_{1}+\alpha_{2}$ is configured to be close to 1, but strict equality is not enforced.
Remark 2.7.
The theoretical analysis of GRPO-S is analogous. See Appendix B.4.
@@
Figure 3: Mean reward trajectories on the test sets. All curves are smoothed for visual clarity. Each row corresponds to a different experimental setting: (Top) AIME 2024 with Qwen2.5-32B, (Middle) AIME 2025 with Qwen2.5-32B, and (Bottom) AIME 2025 with Qwen2.5-7B. Columns show different metrics from left to right: Mean Reward, Mean Reward of Pass@2, Pass@8, and Pass@32. For brevity and to maintain visual clarity, the corresponding results for Pass@4 and Pass@16, which exhibit similar trends, are deferred to Appendix E.3, see Fig. 8 and Fig. 9.
3 Experiments
3.1 Experimental Setup
Tasks and Datasets. Methods are evaluated on the AIME 2024, AIME 2025, and MATH 500 benchmarks (Hendrycks et al., 2021; Lightman et al., 2023). These challenging mathematical datasets require long-horizon thinking(Cobbe et al., 2021), making them a rigorous testbed for assessing advanced reasoning capabilities.
Evaluation Metrics. Our primary metric is Pass@k ($k\in\{2,4,8,16,32\}$), which encourages solution diversity over the more conservative Pass@1. Mean@32 is also reported. Improvements are quantified by Absolute (APG) and Relative (RPG) Performance Gains over the baseline.
Models and Baselines. Experiments are conducted on Qwen2.5-7B/32B (Qwen et al., 2025) models. Comparisons are made against faithful implementations of GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025), as well as the advanced DAPO w/ Forking Tokens (Wang et al., 2025a), to ensure a comprehensive evaluation of baselines.
Implementation Details. Experiments are run on 64 GPUs with a global batch size of 128, a group size of 16, and a learning rate of $1\times 10^{-6}$. For generation, a temperature 1.0 and top-p 1.0 are used, with max lengths of 2048 (prompt) and 4096 (response). Key reward shaping hyperparameters are $\alpha_{1}=\beta_{1}=1$, $\alpha_{2}=\beta_{2}=0.1$, and entropy is clipped at $\epsilon_{low}=0.2,\epsilon_{high}=0.28$. Furthermore, to mitigate length bias, we employ the Group Relative Overlong Punishment mechanism detailed in Appendix D, which adaptively penalizes excessive token lengths based on the classification of problem difficulty (e.g., easy vs. hard).
3.2 Comparative Performance Analysis
As presented in Table 1, our methods, GTPO and GRPO-S, establish a new state-of-the-art by consistently outperforming GRPO, DAPO, and the advanced DAPO w/ Forking Tokens across all configurations, including the rigorous MATH 500 benchmark. Notably, our approach fundamentally differs from DAPO w/ Forking Tokens, which rigidly discards low-entropy tokens via a binary mask. Instead, we employ a continuous entropy-based reward shaping mechanism that fully leverages the learning signal across the distribution. This strategy successfully resolves the stability-exploration trade-off: while DAPO achieves stability at the cost of exploration (plateauing on challenging tasks like AIME 2025), our methods significantly elevate exploration-sensitive scores. This efficacy is most pronounced on smaller models; specifically, regarding the Pass@32 metric on AIME 2024, GTPO achieves a massive absolute performance gain (APG) of +29.4 points on the 7B model, compared to +9.9 points on the 32B model, confirming that entropy-weighting provides a vital learning signal to prevent premature convergence.
3.3 Reward Trajectories and Sample Efficiency